# A Note on Sample Complexity of Interactive Imitation Learning with Log Loss

### v1

@@ 2 Preliminaries
Importantly, $\pi_{u}$ is not a stationary policy; as a result, $a_{1:H}$ are dependent conditioned on $s_{1:H}$; $a_{1:H}$ are only conditionally independent given $s_{1:H}$ and the random policy $\pi$ drawn.
Definition 2 (Each-step mixing of $\mathcal{B}$).
$\bar{\Pi}_{\mathcal{B}}=\{\bar{\pi}_{u,h}(a|s):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a|s):u\in\Delta(\mathcal{B})\}.$
In words, an each-step mixing policy $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ can be executed by drawing $\pi\sim u$ freshly-at-random at each step $h\in[H]$ and plays $\pi_{h}(a|s)$ (e.g. [9, 10]). Observe that $\bar{\pi}_{u}$ is a Markovian policy.
Additionally, we use $\pi(\cdot\parallel s_{1:H})$ to denote the causally-conditioned probability of action sequence $a_{1:H}$ induced by $\pi$, given state sequence $s_{1:H}$ [21].  [^1] The use of $\parallel$ highlights its distinction from standard conditioning on $s_{1:H}$. To elaborate:
@@
It is well-known that the trajectory distribution induced by Markovian policies and their first-step mixtures $\pi$ can be factorized to the product of $\pi(a_{1:H}\parallel s_{1:H})$ and the causally-conditioned probability of the state sequence given the action sequence (Definition 23 and Lemma 24). When it is clear from context, we use shorthand $\pi(s_{1:H})$ for $\pi(\cdot\parallel s_{1:H})$.
Interactive imitation learning protocol.
We consider two oracles that models interacting with demonstration expert: first, the trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ that takes into a state-sequence $s_{1:H}$ and returns $a^{*}_{1:H}\sim\pi^{\mathrm{E}}(\cdot\parallel s_{1:H})$; second, the state-wise demonstration oracle $\mathcal{O}^{\mathrm{State}}$ that takes into a state $s$ and step $h$ and returns $a^{*}\sim\pi^{\mathrm{E}}_{h}(\cdot\parallel s)$.
We consider interactive IL with the learner interacting with the environment and demonstration oracles using the following protocol:
  • For $i=1,2,\dots$
    • Select policy $\pi^{i}$ and rollout in $\mathcal{M}$, observing trajectory $(s_{1},a_{1},\ldots,s_{H},a_{H})$.
    • Query the available oracle(s) to obtain expert annotations.
  • Goal: Return policy $\hat{\pi}$ such that $J(\pi^{\mathrm{E}})-J(\hat{\pi})$ is small, with a few number of queries to $\mathcal{O}^{\mathrm{Traj}}$ or $\mathcal{O}^{\mathrm{State}}$.
By replacing the query of $\mathcal{O}^{\mathrm{Traj}}$ on $s_{1:H}$ with $H$ separate queries to $\mathcal{O}^{\mathrm{State}}$, any algorithm that makes $N$ queries to $\mathcal{O}^{\mathrm{Traj}}$ can be (trivially) converted to one that makes $NH$ queries to $\mathcal{O}^{\mathrm{State}}$. We will later use this observation to make comparison in state-wise annotation model.
Our IL algorithm design will rely on basic online learning algorithms as primitives. An online learning oracle $\mathbb{A}$ with decision space $U$ is said to achieve a regret of $\operatorname{Reg}$ with benchmark set $V$, if it repeatedly chooses $u^{n}\in U$ and then observes loss $\ell^{n}:U\to\mathbb{R}$, and satisfies
$\sum_{n=1}^{N}\ell^{n}(u^{n})-\min_{v\in V}\sum_{n=1}^{N}\ell^{n}(v)\leq\operatorname{Reg}.$
3 Main Results
3.1 Interactive IL enables better sample complexity with state-wise annotation
We first present a interactive IL algorithm, Algorithm 1, that works by interacting with the state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$.
Algorithm 1 DAgger with state-wise annotation oracle
2:  for $n=1,\dots,N$ do
3:     Query $\mathbb{A}$ and receive $u^{n}\in\Delta(\mathcal{B})$.
4:     Execute $\pi^{n}:=\bar{\pi}_{u^{n}}$ and sample $(s^{n},h^{n})\sim d^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{State}}$ for $a^{*,n}=\pi^{\mathrm{E}}_{h^{n}}(s^{n})$.
5:     Update $\mathbb{A}$ with loss function
$\ell^{n}(u):=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a^{*,n}|s^{n})}\right).$ (1)
6:  end for
7:  Output $\hat{\pi}$, a first-step uniform mixture of policies in $\left\{\pi^{1},\ldots,\pi^{n}\right\}$.
Algorithm 1 requires base policy class $\mathcal{B}$ and uses each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ (recall Definition 2). At round $n$, it rolls out the current policy $\pi^{n}=\bar{\pi}_{u^{n}}$, whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples $(s^{n},h^{n})$ from $d^{\pi^{n}}$. The online learning oracle $\mathbb{A}$ it uses has decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$; an classical example of this is the exponential weight algorithm, which has regret bound $\operatorname{Reg}=\ln B$ (see Proposition 21). It then queries $\mathcal{O}^{\mathrm{State}}$ to get $a^{*,n}$ and updates $\mathbb{A}$ by loss $\ell^{n}(u)$ induced by the new example (Eq. (1)). The final policy $\hat{\pi}$ is returned as a uniform first-step mixture of the historical policies $\{\pi^{n}\}_{n=1}^{N}$. We show the following performance guarantee of Algorithm 1:
Theorem 3.
If Algorithm 1 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu H\cdot\frac{\log(B)+2\log(1/\delta)}{N}.$
Theorem 3 shows that Algorithm 1 returns a policy of suboptimality $O(\frac{\mu H\ln B}{N})$ using $N$ state-wise annotations from the expert. In comparison, with $N$ state-wise annotations, one can obtain $N/H$ trajectory-wise annotations; [4]’s analysis shows that behavior cloning on $N/H$ iid trajectories from $\pi^{\mathrm{E}}$ returns a policy of suboptimality $O(\frac{RH\ln B}{N})$. As $\mu$ is no greater than $R$ and can be much smaller, Algorithm 1 has a faster convergence rate than behavior cloning under the same state-wise annotation budget.
[…]
$\operatorname{OnEst}^{\mathrm{State}}_{N}:=\sum_{n=1}^{N}\mathbb{E}_{(s,h)\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right].$
The proof involves bounding the state-wise Hellinger estimation error and translating it into the performance difference difference beteween $\hat{\pi}$ and $\pi^{\mathrm{E}}$. More details are provided in Appendix C.
3.2 Interactive IL matches offline IL on trajectory-wise annotation
@@
Next, we consider the trajectory-wise sampling model. We present another DAgger variant, namely Algorithm 2 and provide its sample complexity bounds.
@@
Algorithm 2 DAgger with trajectory-wise annotation oracle
@@
$\ell^{n}(u):=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a_{1:H}^{*,n}\parallel s_{1:H}^{n})}\right).$ (2)
@@
Algorithm 2 uses first-step mixing policies $\pi_{u}\in\Pi_{\mathcal{B}}$ (recall Definition 1). At round $n$, it rolls out $\pi^{n}=\pi_{u^{n}}$ whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples a full state sequence $s^{n}_{1:H}$. Same as Algorithm 1, Algorithm 2 also requires $\mathbb{A}$ to have decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$. It then requests expert’s trajectory-wise annotation $a^{*,n}_{1:H}$ and updates $\mathbb{A}$ by $\ell^{n}(u)$ (Eq. (2)). At the end of round $N$, the uniform first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ is returned, which is equivalent to returning $\pi_{\hat{u}}$, where $\hat{u}:=\frac{1}{H}\sum_{n=1}^{N}u^{n}$. We provide the following performance guarantee of Algorithm 2:
@@
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
Theorem 4 shows that the interactive IL Algorithm 2 matches the trajectory-wise sample complexity of behavior cloning in [4]. In contrast, prior state-of-the-art analysis of interactive IL algorithms [4, Appendix C.2] gives sample complexity results that are in general worse than behavior cloning.  [^2] For [4, Appendix C.2]’s sample complexity to improve over behavior cloning, we need $\mu H\max_{h\in[H]}\ln|\mathcal{B}_{h}|$ to be significantly smaller $R\ln|\mathcal{B}|$ (where $\mathcal{B}_{h}$ is the projection of $\mathcal{B}$ onto step $h$). This may require the strong condition that $\mu<R/H<1$ in the practically-popular parameter sharing settings ($|\mathcal{B}_{h}|=|\mathcal{B}|$).
@@
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}:=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s_{1:H}),\pi^{\mathrm{E}}(s_{1:H}))\right].$
@@
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}$ decouples the dependency between the state sequence and the distribution of action sequence induced by the learner. Perhaps surprisingly, it is compatible with non-Markovian first-step mixture of policies, while still being well-behaved enough to be translated to a policy suboptimality guarantee, which could be of independent interest. Further details can be found in Appendix B.
3.3 Summary of Results
Table 1: Sample complexity of imitation learning algorithms to achieve $J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\epsilon$. In the table, the cumulative cost satisfies $R\leq H$, and the recoverability factor is bounded by $1\leq\mu\leq H$.
Algorithms Annotated Trajectories Annotated States
LogLossBC [4] $O\left(\frac{R\log(B)}{\epsilon}\right)$ $O\left(\frac{RH\log(B)}{\epsilon}\right)$
LogLossDAgger [4] $O\left(\frac{\mu H\log(B)}{\epsilon}\right)$ $O\left(\frac{\mu H^{2}\log(B)}{\epsilon}\right)$
Algorithm 1 (Ours) $O\left(\frac{\mu H\log(B)}{\epsilon}\right)$
Algorithm 2 (Ours) $O\left(\frac{R\log(B)}{\epsilon}\right)$ $O\left(\frac{RH\log(B)}{\epsilon}\right)$
Table 1 summarizes the number of annotations needed to achieve $J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\epsilon$ for both state-wise and trajectory-wise annotation settings. We include baselines from recent advancements in imitation learning, covering both offline IL (LogLossBC [4]) and interactive IL (LogLossDAgger [4]).  [4] originally does not provide sample complexity results for the state-wise annotation setting; to enable direct comparison, we convert their trajectory-wise sample complexities to state-wise sample complexities by multiplying by $H$.
For the state-wise annotation setting, DAgger with state-wise annotation oracle demonstrates significant improvements over all baselines when $\mu\ll R$. For the trajectory-wise annotation setting, DAgger with trajectory-wise annotation oracle achieves horizon-independent guarantees without the $\mu$-recoverability assumption, matching the minimax optimal performance of LogLossBC [4].
These findings provide two key clarifications: 1. Interactive IL offers potential advantage over offline IL in terms of state-wise sampling efficiency. 2. Interactive IL does not inherently suffer from suboptimal sample complexity in general compared to offline IL.
4 Related Works and Discussions
State-wise v.s. trajectory-wise annotations in IL.
The debate over using state-wise versus trajectory-wise annotations traces back to the reduction from imitation learning to supervised learning [13]. While the sample complexity of BC can be interpreted in a state-wise manner (e.g. [1, 9]), BC conventionally relies on full trajectories of expert demonstrations. Recent advancements [4] seem to settle the debate in favor of trajectory-wise annotations through more refined analysis. However, the question ”Is trajectory-wise annotation all you need?” remains unresolved in the interactive setting.
Unlike offline demonstrations, interactive annotations require querying effort. As noted by [11], “A 30-second manipulation task with 20 Hz robot control would require a human to relabel 600 state samples.” When trajectory annotation costs scale with the number of steps, annotating a single sampled state can be far more cost efficient.
Though the use of state-wise annotation appeared early in the interactive IL literature, such as in the Forward Training algorithm [13], its adoption [6] was less influential compared to DAgger [15], the interactive counterpart of BC, which popularized the convention of annotating full state sequences. Subsequent works following DAgger diverged into state-wise [9, 19, 10], trajectory-wise [20, 16, 4], and other variants [14, 1].To the best of our knowledge, we are the first to explicitly address this debate, systematically developing efficient algorithms tailored to both state-wise and trajectory-wise annotations.
First-step mixing and every-step mixing policies.
The emergence of first-step mixing originated from technical considerations. In may interactive IL methods [15, 14, 14], the returned policy was not a uniform first-step mixture but rather the best policy selected through validation. However, performing such validation in an interactive setting often requires additional expert annotations. Subsequent works [16, 9, 10, 4] circumvented the need for validation by employing a uniform first-step mixture of policies across learning rounds, thereby directly translating online regret guarantees into performance differences. While [20] implies the usage of a first-step mixture policy class, to the best of our knowledge, we are the first to establish a concrete online regret guarantee this policy class.
On the other hand, every-step mixing between the learned policy across rounds and the expert policy has been a prevalent strategy in interactive IL approaches [3, 13, 15, 14, 20]. However, when selecting the mixture policy class for each learning round, ambiguity arises regarding the choice between first-step mixing and every-step mixing [20]. For every-step mixture policies, [9] was the first to explicitly distinguish this approach from first-step mixing. In other works [16, 4], every-step mixing can be interpreted as learning $H$ separate mixture policies, one for each step within an episode.
Alternative algorithm designs and practical implementations.
Though this work follows [4] and focuses on log loss, we believe the $1/n$ rate is not exclusive to log loss. Despite requiring an additional supervision oracle, [8] suggests that trajectory-wise annotation complexity similar to Theorem 4 can be achieved using Halving [18] and 00-$1$ loss.
From an algorithmic perspective, we explored trajectory-wise annotation with first-step mixing (Algorithm 2) and state-wise annotation with every-step mixing (Algorithm 1). For trajectory-wise annotation with every-step mixing, naively learning a parameter-sharing policy may encounter a batch-summed log loss, introducing an additional $H$ factor to the sample complexity, which is undesirable. State-wise annotation with every-step mixing remains an open question for future research.
In terms of practical implementation, it is worth noting that even with oracle-efficient implementations (e.g. [9, 10]), interactive IL requires multiple computational oracle calls per learning round. In contrast, offline IL requires only a single oracle call to obtain the returned policy, which provides a clear computational advantage.
Lower bounds in interactive IL.
From an information-theoretic perspective, a line of work [17, 16, 4] provides lower bounds for imitation learning under the realizable setting and considers $\mu$-recoverability. [16] is the first to demonstrate a gap between the lower bounds of offline IL and interactive IL in trajectory-wise annotation, focusing on the non-parameter-sharing setting. [4] shifts attention back to the general setting and establishes a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for trajectory-wise annotation.
We argue that the result in [4] also implies a $\Omega(\frac{H}{\epsilon})$ sample complexity lower bound for the state-wise annotation setting. Their proof relies on an MDP consisting only of self-absorbing states, where annotating a full trajectory is equivalent to annotating a single state. In their special MDP case ($1$-recoverable), Algorithm 1 achieves $\tilde{O}(\frac{H\log(B)}{\epsilon})$ state-wise sample complexity, which does not contradict existing lower bounds. Nonetheless, obtaining lower bounds for state-wise sample complexity for general MDPs and policy classes remains an open question.
Conclusion.
We present new results of interactive IL with log loss that matches the performance of offline IL in trajectory-wise annotation and achieves improved sample complexity in state-wise annotation. Our analysis introduced a new notion of decoupled Hellinger estimation error, a powerful tool for analyzing interactive IL algorithms. We leave the exploration of stochastic or non-realizable expert settings open for future investigation.
Appendix A Useful Distance Measures
In the following, we present 4 useful distance measures for pair of policies.
Definition 5 (Trajectory-wise $L_{1}$-divergence).
$\lambda(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(s_{1:H})}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq a^{\prime}_{h})\right].$
$\lambda(\pi\parallel\pi^{\prime})$ is the expected total number of actions taken by $\pi^{\prime}$ that deviates from actions in trajectories induced by $\pi$. Note that $\lambda(\cdot||\cdot)$ is asymmetric, while the same concept is applied in offline and interactive IL [13, 15] with different guarantees for $\lambda(\pi\parallel\pi^{\mathrm{E}})$ and $\lambda(\pi^{\mathrm{E}}\parallel\pi)$ (Lemma 19).
Definition 6 (Trajectory-wise $L_{\infty}$-semi-metric [4]).
$\rho(\pi\parallel\pi^{\prime})$ is the probability of any action taken by $\pi^{\prime}$ deviating from actions in trajectories induced by $\pi$, which is symmetric [4]. A bound on $\rho(\pi\parallel\pi^{\mathrm{E}})$ leads to straightforward performance difference guarantee: $J(\pi)-J(\pi^{\mathrm{E}})\leq R\cdot\rho(\pi\parallel\pi^{\mathrm{E}})$ [4] (Lemma 20).
Definition 7 (State-wise Hellinger distance).
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their state-wise Hellinger distance as $\mathbb{E}_{(s,h)\sim d^{\pi}}\left[D^{2}_{\text{H}}(\pi_{h}(s),\pi^{\prime}_{h}(s))\right]$.
State-wise Hellinger distance represents the expected Hellinger distance between the action distribution of $\pi$ and $\pi^{\prime}$ on $(s,h)\sim d^{\pi}$. One notable feature here is that the distance is evaluated between $\pi_{h}(\cdot\mid s)$ and $\pi^{\prime}_{h}(\cdot\mid s)$, independent of the original action $a$ taken by $\pi$ when visiting $s$. By Lemma 19, state-wise Hellinger distance can relate to trajectory-wise $L_{1}$-divergence.
Definition 8 (decoupled Hellinger distance).
Appendix B Decoupling States and Actions: New Guarantees for DAgger
B.1 Decoupling State and Action Sequences by Decoupled Hellinger Distance

### v2

@@ E.1 Additional Notations and Useful Distance Measures
Importantly, $\pi_{u}$ is not a stationary policy (note its difference with $\bar{\pi}_{u}$ in Definition 1); as a result, $a_{1:H}$ are dependent conditioned on $s_{1:H}$, while $a_{1:H}$ are only conditionally independent given $s_{1:H}$ and the random policy $\pi$ drawn.
Additionally, we use $\pi(\cdot\parallel s_{1:H})$ to denote the causally-conditioned probability of action sequence $a_{1:H}$ induced by $\pi$, given state sequence $s_{1:H}$ [ziebart2010modeling].  [^2] The use of $\parallel$ highlights its distinction from standard conditioning on $s_{1:H}$. To elaborate:
@@
It is well-known that the trajectory distribution induced by Markovian policies and their first-step mixings $\pi$ can be factorized to the product of $\pi(a_{1:H}\parallel s_{1:H})$ and the causally-conditioned probability of the state sequence given the action sequence (Definition 40 and Lemma 41). When it is clear from context, we use shorthand $\pi(s_{1:H})$ for $\pi(\cdot\parallel s_{1:H})$.
In the following, we present another 2 useful distance measures for pair of policies.
Definition 25 (Trajectory-wise $L_{\infty}$-semi-metric [foster2024behavior]).
$\rho(\pi\parallel\pi^{\prime})$ is the probability of any action taken by $\pi^{\prime}$ deviating from actions in trajectories induced by $\pi$, which is symmetric [foster2024behavior]. A bound on $\rho(\pi\parallel\pi^{\mathrm{E}})$ leads to straightforward performance difference guarantee: $J(\pi^{\mathrm{E}})-J(\pi)\leq R\cdot\rho(\pi\parallel\pi^{\mathrm{E}})$ [foster2024behavior] (Lemma 37).
Definition 26 (Decoupled Hellinger distance).
E.2 Interactive IL Matches Offline IL on Trajectory-wise Annotation
@@
Next, we consider the trajectory-wise sampling model. We present Tragger, another DAgger variant, namely Algorithm 3 and provide its sample complexity bounds.
@@
Algorithm 3 Tragger: DAgger with trajectory-wise annotation oracle
@@
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi^{n}(a_{1:H}^{*,n}\parallel s_{1:H}^{n})}\right).$ (5)
@@
Algorithm 3 uses first-step mixing policies $\pi_{u}\in\Pi_{\mathcal{B}}$ (recall Definition 24). At round $n$, it rolls out $\pi^{n}=\pi_{u^{n}}$ whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples a full state sequence $s^{n}_{1:H}$. Same as Algorithm 1, Algorithm 3 also requires $\mathbb{A}$ to have decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$. It then requests expert’s trajectory-wise annotation $a^{*,n}_{1:H}$ and updates $\mathbb{A}$ by $\ell^{n}(\pi)$ (Eq. (5)). At the end of round $N$, the uniform first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ is returned, which is equivalent to returning $\pi_{\hat{u}}$, where $\hat{u}:=\frac{1}{H}\sum_{n=1}^{N}u^{n}$. We provide the following performance guarantee of Algorithm 3:
@@
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq 2R\frac{\log(B)+2\log(1/\delta)}{N}.$
Theorem 27 shows that the interactive IL Algorithm 3 matches the trajectory-wise sample complexity of behavior cloning in [foster2024behavior]. In contrast, prior state-of-the-art analysis of interactive IL algorithms [foster2024behavior, Appendix C.2] gives sample complexity results that are in general worse than behavior cloning.  [^3] For [foster2024behavior, Appendix C.2]’s sample complexity to improve over behavior cloning, we need $\mu H\max_{h\in[H]}\log|\mathcal{B}_{h}|$ to be significantly smaller $R\log|\mathcal{B}|$ (where $\mathcal{B}_{h}$ is the projection of $\mathcal{B}$ onto step $h$). This may require the strong condition that $\mu<R/H<1$ in the practically-popular parameter sharing settings ($|\mathcal{B}_{h}|=|\mathcal{B}|$).
@@
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}:=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
@@
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}$ decouples the dependence between the state sequence and the distribution of action sequence induced by the learner. Perhaps surprisingly, it is compatible with non-Markovian first-step mixing of policies, while still being well-behaved enough to be translated to a policy suboptimality guarantee, which could be of independent interest.
E.3 Decoupling State and Action Sequences by Decoupled Hellinger Distance